# Arithmetic on Continued Fractions

### v1

@@ 3.5. Computing the range of $M$
$\left\{\frac{a^{\prime}}{e^{\prime}},\frac{b^{\prime}}{f^{\prime}},\frac{c^{\prime}}{g^{\prime}},\frac{d^{\prime}}{h^{\prime}}\right\}\ .$
4. Solving Quadratic Equations
4.1. Square root of a rational number
The arithmetic algorithm can be extended to solve quadratic equations with CF coefficients. We begin with the special case of computing the square root of a rational number $\alpha$. The key idea is that $\sqrt{\alpha}$ is the fixed point of a homographic function, i.e. the solution to
$y=\frac{\alpha}{y}\ .$
We thus need to find a continued fraction $y$ which produces $y$ given the initial matrix $\left(\begin{smallmatrix}0&p\\ q&0\end{smallmatrix}\right)$, where $\alpha=\frac{p}{q}$.
As a concrete example, we find $\sqrt{11}$. The first term will be the floor of the fixed point of $M=\left(\begin{smallmatrix}0&11\\ 1&0\end{smallmatrix}\right)$. Of course it is easy to observe in this case that $M(3)>3$ and $M(4)<4$, so the first term is $y_{0}=3$; but we must find a better algorithm than linear search through $M(1),M(2),M(3)\cdots$. Such an algorithm exists because any homographic function of the form
$\begin{pmatrix}p&q\\ r&-p\end{pmatrix}$
is self-inverse: i.e. $M(M(y))=y$. This immediately implies that, given any “initial guess” $\hat{y}$, the fixed point is between $\hat{y}$ and $M(\hat{y})$. So we can use binary search, starting perhaps from $\hat{y}=1$, to find the desired term in logarithmic time. When $\hat{y}=\lfloor M(\hat{y})\rfloor$ or $\hat{y}=\lceil M(\hat{y})\rceil$, this common value is the next term of $y$. The self-inverse property is preserved by ingesting and then producing the same term:
$\mbox{produce}\left(\hat{y},\mbox{ingest}\left(\hat{y},\left(\begin{smallmatrix}p&q\\ r&-p\end{smallmatrix}\right)\right)\right)=\left(\begin{smallmatrix}\hat{y}r-p&r\\ q+2\hat{y}p-\hat{y}^{2}r&p-\hat{y}r\end{smallmatrix}\right)\ .$
Thus each subsequent iteration can be handled the same way. After $y_{0}=3$ we obtain
$M=\begin{pmatrix}3&1\\ 2&-3\end{pmatrix}$
which yields $y_{1}=3$ since now $M(3)=\frac{10}{3}$. Ingesting and producing 3 leads to
$M=\begin{pmatrix}3&2\\ 1&-3\end{pmatrix}$
implying $y_{2}=6$ because $M(6)=\frac{20}{3}$. The next iteration puts us in a loop: ingesting and producing 6 takes us back to
$\begin{pmatrix}3&1\\ 2&-3\end{pmatrix}$
proving that $\sqrt{11}=[3,(3,6)]$.
4.2. Square root of an explicit CF
Now we consider how to extract the square root of a real number $x$ given as a continued fraction. We will start with the simpler case in which $x$ expressed as an explicit list of integer terms, before generalizing to an algorithm that operates on our internal representation of a CF as a list of intervals.
The square root of $x$ is the fixed point $y=\frac{x}{y}$. The matrix $\left(\begin{smallmatrix}0&1&0&0\\ 0&0&1&0\end{smallmatrix}\right)$ is bihomographic, but we can also think of it as a homographic matrix $\left(\begin{smallmatrix}0&x\\ 1&0\end{smallmatrix}\right)$ with constant $x$ and variable $y$. More generally let $M_{\mathbf{x}}$ denote $M$ with $x$ fixed at $\mathbf{x}$, i.e. if $M=\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right)$ then
$M_{\mathbf{x}}=\begin{pmatrix}(a\mathbf{x}+c)&(b\mathbf{x}+d)\\ (e\mathbf{x}+g)&(f\mathbf{x}+h)\end{pmatrix}\ .$
As before, we use boldface $\mathbf{x}$ to distinguish a specific input from the symbolic variable. $M_{\infty}$ is the limit $\left(\begin{smallmatrix}a&b\\ e&f\end{smallmatrix}\right)$.
Starting from $\left(\begin{smallmatrix}0&1&0&0\\ 0&0&1&0\end{smallmatrix}\right)$, we must ingest terms of $\mathbf{x}$ until we reach a stage at which the floor of the fixed point does not depend on $x$; in other words, until the floor of the fixpoint is the same when $x=1$ and $x=\infty$. This floor will be the next term of $y$. To find the next term, we first use the same binary search as before to get $\hat{y}$, the floor of the fixed point of of $M_{1}$; then we check whether $\hat{y}$ equals $\lfloor M_{\infty}(\hat{y})\rfloor$ or $\lceil M_{\infty}(\hat{y})\rceil$. If so, we ingest $\hat{y}$ and produce $\hat{y}$; if not, we ingest another term of $\mathbf{x}$ and continue.
This works because the matrices $M_{x}$ will always have the self-inverse property. Such is clearly the case with the initial matrix, and a similar computation to what we did for rational $\mathbf{x}$ proves the property is preserved when we ingest and produce a term.
As an example, we compute the fourth root of 2, i.e. the square root of $\mathbf{x}=[1,(2)]$. Making the substitution $x\leftarrow 1+\frac{1}{x}$ gives
$M=\frac{x+1}{xy}\ .$
The fact that $\lfloor\mathbf{x}\rfloor=1$ should be enough to imply that $\lfloor y\rfloor=1$, and indeed this is the case:
$\begin{array}[]{cc}M_{1}=\frac{2}{y}&M_{\infty}=\frac{1}{y}\\ M_{1}(1)=2&M_{\infty}(1)=1\end{array}$
Ingesting and producing $y_{0}=1$ leads to $M=\frac{yx+x}{y-x}$; then ingesting 2 from $\mathbf{x}$ gives
$M=\frac{2xy+2x+y+1}{xy-2x-1}\ .$
We do not yet have enough information to determine the next output term:
$\begin{array}[]{cc}M_{1}=\frac{3y+3}{y-3}&M_{\infty}=\frac{2y+2}{y-2}\\ M_{1}(6)=7&M_{\infty}(5)=4\end{array}$
Ingesting another 2 from $\mathbf{x}$ will suffice:
$M=\frac{5xy+5x+2y+2}{2xy-5x+y-2}$
@@ 4.2. Square root of an explicit CF
$\begin{array}[]{cc}M_{1}=\frac{7y+7}{3y-7}&M_{\infty}=\frac{5y+5}{2y-5}\\ M_{1}(5)=\frac{21}{4}&M_{\infty}(5)=5\end{array}$
so $y_{1}=5$. After ingesting/producing 5 and then reading another 2 from $\mathbf{x}$ we reach
$M=\frac{13xy+5x+5y+2}{7xy-13x+5y-5}$
which tells us that $y_{2}=3$: $M_{1}(3)=\frac{61}{18}$ and $M_{\infty}(4)=\frac{19}{5}$. The first few terms of $4]{2}$ are $1.18921\cdots=[1,5,3,1,1,40,5,\cdots]$.
  {input $x$: CF represented as list of intervals}
  {output $y$: CF expansion of $\sqrt{x}$ represented as list of intervals}
  $i\leftarrow 0$index of the next element of $x$ we will read}
  $j\leftarrow 0$index of the next element of $y$ we will generate}
  $(x_{\ell},x_{h})\leftarrow(1,\infty)$current bounds on remaining part of $x$}
  $M=\left(\begin{smallmatrix}0&1&0&0\\ 0&0&1&0\end{smallmatrix}\right)$solve $y=\frac{x}{y}$}
  while $M\neq\infty$ do
     {read input}
     if $x_{i}=[a,b)$ is ambiguous then
        $(x_{\ell},x_{h})\leftarrow(a,b)$
     else
        $M\leftarrow\mbox{ingest\_x}(x_{i},M)$
        $(x_{\ell},x_{h})\leftarrow[1,\infty)$
     end if
     $i\leftarrow i+1$
     {produce output if we can}
     $\hat{y}\leftarrow$ floor of fixed point of $M_{x_{\ell}}$found with binary search}
     if floor of fixed point of $M_{x_{h}}$ equals $\hat{y}$ then
        $y_{j}\leftarrow\hat{y}$
        $j\leftarrow j+1$
        $M\leftarrow\mbox{produce}(\hat{y},\mbox{ingest\_y}(\hat{y},M))$
     end if
  end while
Figure 4. Algorithm for square root of a CF
4.3. Quadratic equations
A quadratic equation can be written as the fixed point of a self-inverse homographic function: $py^{2}+qy+r=0$ is equivalent to
(13) $y=\frac{-qy-2r}{2py+q}\ .$
If $q$ is a continued fraction, we can solve (13) by starting with $M=\left(\begin{smallmatrix}-x&-2r\\ 2p&x\end{smallmatrix}\right)$ and finding the fixed point, ingesting terms of $q=\mathbf{x}$. The cases of $p$ or $r$ as continued fractions are similar.
4.4. Square root of a general CF
The algorithm is not much different for $x$ represented as a list of intervals. When we read an ambiguous interval $[x_{\ell},x_{h})$, we do not change the current matrix; instead we check whether the floors of the fixed points of $M_{x_{\ell}}$ and $M_{x_{h}}$ are the same. If so, this common value is the next term of $y$.
Note we have assumed that all terms of $x$, including the first term, are greater than or equal to 1. If $x=[0,x_{1},x_{2},\cdots]$ then $\sqrt{x}=[0,z_{0},z_{1},\cdots]$ where $z=\sqrt{[x_{1},x_{2},\cdots]}=\sqrt{1/x}$.

### v2

@@ 3.4 Computing the range of $M$
$\left\{\frac{a^{\prime}}{e^{\prime}},\frac{b^{\prime}}{f^{\prime}},\frac{c^{\prime}}{g^{\prime}},\frac{d^{\prime}}{h^{\prime}}\right\}$
assuming no sign change in the denominator as $x^{\prime}$ and $y^{\prime}$ go from 0 to infinity; fractions with numerator and denominator both 0 can be ignored. If the denominator can be zero, $\rho\left(\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right),[x_{\ell},x_{u}),[y_{\ell},y_{u})\right)=[-\infty,\infty)$
3.5 Possible Optimization
When we have to read the next interval from $x$ or $y$, and we see an ambiguous interval at the head of the list, it is not necessary to use that interval; we may skip it and go on to the next element of the list, which will give a tighter bound on the next term. Any such heuristic must guarantee that, given an infinite list of ambiguous input intervals, we eventually read something and recompute the range of $M$. Similarly, the algorithm does not need to output every ambiguous interval of $z$; but we must guarantee that we eventually output something, even if we are generating an infinite list of ambiguous intervals. The most obvious idea is to read and generate ambiguous intervals only when the width of the interval is smaller than some threshold.
3.6 Division by Zero
The limit of $\frac{axy+bx+cy+d}{exy+fx+gy+h}$ as $x$ or $y$ approaches infinity is not necessarily a homographic matrix; for example
$\lim_{y\to\infty}\frac{y}{x-3}=\begin{cases}+\infty&\mbox{if }x>3\\ \mbox{undefined}&\mbox{if }x=3\\ -\infty&\mbox{if }x<3\end{cases}$
But ingesting infinity can only mean that we have reached the end of a finite rational input $\mathbf{y}$. At a previous step we must have ingested an explicit integer term $\alpha$. The previous matrix, which generated $\frac{y}{x-3}$ by substituting $y\leftarrow\alpha+1/y$, can be found by applying the inverse substitution $y\leftarrow 1/(y-\alpha)$ to get
$\frac{1}{(y-\alpha)(x-3)}=\frac{1}{xy-\alpha x-3y+3\alpha}$
which means we are trying to divide by zero; we are substituting $y=\alpha+\frac{1}{\infty}=\alpha$. The invalid computation is revealed by the anomalous limit. Such anomalous limits can only arise from division by zero.
4 Transcendental Functions
4.1 Exponentials
We now consider how to compute the CF terms of
$e^{x}=\sum_{k=0}^{\infty}\frac{x^{k}}{k!}$
where $x$ is given as a CF, possibly with ambiguous intervals. Let
$M_{n,x}(y)=1+\frac{xy}{n}\ .$
This is a bihomographic expression with matrix $\left(\begin{smallmatrix}1&0&0&n\\ 0&0&0&n\end{smallmatrix}\right)$. The $n^{th}$ degree Taylor approximation, $1+x+x^{2}/2+\cdots+x^{n}/{n!}$, can be written using Horner’s rule as
$1+x\left(1+\frac{x}{2}\left(1+\frac{x}{3}\left(1+\cdots\left(1+\frac{x}{n}\right)\right)\right)\cdots\right)=M_{(1,x)}(M_{(2,x)}(\cdots M_{(n,x)}(1)\cdots))$
so in the limit
[…]
Now define
$y_{n}(x)=1+\frac{x}{n}+\frac{x^{2}}{n(n+1)}+\cdots=\lim_{m\to\infty}M_{(n,x)}(M_{(n+1,x)}(\cdots M_{(m,x)}(1)\cdots))\ .$ (8)
Thus $y_{n}(x)=M_{(n,x)}(y_{n+1}(x))$, and $y_{1}(x)=e^{x}$. Assume that $x\geq 0$ (otherwise we can use the identity $e^{-x}=1/e^{x}$); we have
$y_{n}(x)-1\leq\frac{e^{x}-1}{n}\ .$
We can always obtain arbitrarily tight rational bounds $x_{\ell}\leq x\leq x_{h}$ from a finite prefix of the CF representation of $x$; since $e<3$ we have the rational bounds
$1+\frac{x_{\ell}}{n}\leq y_{n}(x)\leq 1+\frac{3^{\lceil x_{h}\rceil}-1}{n}\ .$ (9)
If this interval is ambiguous, it can be the first interval of the CF representation of $y_{n}(x)$. The subsequent intervals are the CF representation of $M_{(n,x)}(y_{n+1}(x))$, which is a bihomographic expression that can be evaluated with the arithmetic algorithm; the algorithm will read
$\left(1+\frac{x_{\ell}}{n+1},1+\frac{3^{\lceil x_{h}\rceil}-1}{n+1}\right)$
as the first interval of $y_{n+1}(x)$ and so on. The infinite recursion of (8) does not lead to an infinite loop, because we can always get the first interval of $y_{n}(x)$ without recursion into $y_{n+1}(x)$.
There is one further point to consider; the first interval of $M_{(n,x)}(y_{n+1}(x))$ might not be a subset of (9). Since both intervals contain $y_{n}(x)$, we can take their intersection as the second interval of $y_{n}(x)$. We might need to continue taking such intersections as long as $M_{(n,x)}(y_{n+1}(x))$ produces ambiguous terms.
If the interval (9) is unambiguous, giving $a$ as the first term of $y_{n}(x)$, we can immediately produce $a$, followed by applying the arithmetic algorithm to
$\frac{1}{M_{(n,x)}(y_{n+1}(x))-a}\ .$
4.1.1 Validity of the Algorithm
At this point we have given an algorithm which produces a sequence of intervals for each $y_{n}(x)$; it remains to prove that the sequence of intervals representing $y_{1}(x)$ is in fact valid as defined in section 3.1. The properties of the arithmetic algorithm only guarantee that, if $x$ and $y_{n}(x)$ are valid sequences of intervals, then $y_{n-1}(x)$ is valid.
The algorithm already guarantees that successive ambiguous intervals will be nested; we must prove that the upper and lower bounds converge to the same point. In other words we must prove that, for any $\varepsilon>0$, we eventually obtain a prefix $I_{0},I_{1},\cdots I_{k}]$ of $y_{1}(x)$ which confines $y_{1}(x)$ within an interval of width $\leq\varepsilon$; we say that $y_{1}(x)$ is $\varepsilon$-bounded.
So suppose that at some point we have ingested enough of $y_{n+1}(x)$ to have $y_{n+1}(x)$ be $\varepsilon^{\prime}$-bounded, with $y_{\ell}\leq y_{n+1}(x)\leq y_{\ell}+\varepsilon^{\prime}$. The definition of $M_{(n,x)}$, along with $x_{\ell}\leq x\leq x_{h}$, implies that $y_{n}(x)$ is $\varepsilon$-bounded if
$\varepsilon^{\prime}\leq\frac{n}{x_{h}}\varepsilon-\frac{x_{h}-x_{\ell}}{x_{h}}y_{\ell}$
But $\frac{x_{h}-x_{\ell}}{x_{h}}$ approaches zero as we ingest more of $x$, while $y_{\ell}$ is bounded above by a constant, so eventually it will suffice to have
$\varepsilon^{\prime}\leq\frac{n}{2x_{h}}\varepsilon$
Thus $y_{1}(x)$ is $\varepsilon$-bounded if some $y_{n}(x)$ is $\varepsilon^{\prime}$-bounded with
$\varepsilon^{\prime}\leq\frac{(n-1)!}{(2x_{h})^{n-1}}\varepsilon\ .$
Now $\frac{(n-1)!}{(2x_{h})^{n-1}}$ approaches infinity, while our algorithm immediately bounds each $y_{n}(x)$ within an interval of width
$\frac{3^{\lceil x_{h}\rceil}-1-x_{\ell}}{n}$
which approaches zero; therefore some $y_{n}(x)$ will be $\varepsilon^{\prime}$-bounded as required. The CF will converge to the correct value $e^{x}$, since all intervals bounding $y_{1}(x)$ by construction contain $e^{x}$.
4.2 Logarithms
We can apply the same approach to $\log(x)$, using the series
$\log(x)=2\left(\left(\frac{x-1}{x+1}\right)+\frac{1}{3}\left(\frac{x-1}{x+1}\right)^{3}+\frac{1}{5}\left(\frac{x-1}{x+1}\right)^{5}\cdots+\right)$
which converges for all $x>0$. Let $z=\frac{x-1}{x+1}$, and define
[…]
Then $\log(x)=2zg(z^{2})$, so we only need consider how to compute $g(w)$ with $0\leq w\leq 1$. The series for $g(w)$ can be obtained by iterating
$M_{(n,w)}(y)=1+\frac{2n-1}{2n+1}wy$
i.e. the bihomographic expression $\left(\begin{smallmatrix}2n-1&0&0&2n+1\\ 0&0&0&2n+1\end{smallmatrix}\right)$. So we define
$g_{n}(w)=1+\sum_{k=1}^{\infty}\frac{2n-1}{2n+2k-1}w^{k}=\lim_{m\to\infty}M_{(n,w)}(M_{(n+1,w)}(\cdots M_{(m,w)}(1)\cdots))\ .$
In particular $g_{1}(w)=g(w)$, and $g_{n}(w)=M_{(n,w)}(g_{n+1}(w))$. Since $w$ is nonnegative, we have
$1\leq g_{n}(w)\leq 1+\frac{2n-1}{2n+1}\sum_{k=1}^{\infty}w^{k}=1+\frac{2n-1}{2n+1}\frac{w}{1-w}\ .$ (10)
A finite prefix of the representation of $x$ will suffice to obtain a rational upper bound $B_{x}$ on $\frac{w}{1-w}$, so the first interval of $g_{n}(w)$ can be $(1,1+\frac{2n-1}{2n+1}B_{x})$, followed by applying the arithmetic algorithm to $M_{(n,w)}(g_{n+1}(w))$. However, if $\frac{2n-1}{2n+1}B_{x}\leq 1$, we can immediately produce 1 as the first term, followed by applying the arithmetic algorithm to
$\frac{1}{M_{(n,w)}(g_{n+1}(w))-1}\ .$
4.2.1 Validity
Proceeding along the same lines as section (4.1.1), $g_{n}(w)$ will be $\varepsilon$-bounded if $g_{n+1}(w)$ is $\frac{2n+1}{(2n-1)\hat{w}}\varepsilon$-bounded for some $w<\hat{w}<1$; so $g_{1}(w)$ will be $\varepsilon$-bounded if some $g_{n}(w)$ is $\frac{2n+3}{\hat{w}^{n-1}}\varepsilon$-bounded. This width approaches infinity, while the algorithm immediately confines each $g_{n}(w)$ in an interval no wider than the constant $B_{x}$.
4.3 Trigonometric Functions
The cosine
[…]
can be represented by letting $w=x^{2}$, defining
$M_{n,w}(y)=1-\frac{wy}{2n(2n-1)}$
@@ 4.3 Trigonometric Functions
$c_{n}(w)=M_{n,w}(M_{n+1,w}(\cdots))=1-\frac{w}{2n(2n-1)}+\frac{w^{2}}{(2n+2)(2n+1)2n(2n-1)}-\cdots\ .$ (11)
so $c_{n}(w)=M_{n,w}(c_{n+1}(w))$, and $c_{1}(w)=\cos(x)$.
If $|x|\leq\frac{\pi}{2}$, then $w<\frac{5}{2}$, and (11) is an alternating series bounded between $1-\frac{5}{4n(2n-1)}$ and 1; the CF expansion of $c_{1}(w)$ will start with this interval, followed by the result of applying the arithmetic algorithm to the bihomographic expression $M_{n,w}(c_{2}(w))$. For $n\geq 2$, we can immediately produce 0, and we have
$1\leq\frac{1}{M_{n,w}(c_{n+1}(w))}\leq\frac{4n(2n-1)}{4n(2n-1)-5}\ .$
This interval in turn lets us immediately produce 1; then
$\frac{4n(2n-1)-5}{5}\leq\left(\frac{1}{M_{n,w}(c_{n+1}(w))}-1\right)^{-1}<\infty$
and subsequent intervals come from applying the arithmetic algorithm to
$\left(\frac{1}{M_{n,w}(c_{n+1}(w))}-1\right)^{-1}\ .$
For $\frac{\pi}{2}\leq|x|\leq\pi$ we use $\cos(x)=-\cos(\pi-x)$, and for all other values $\cos(x)=\cos(x\pm 2\pi)$; see section 4.5 for computing the CF expansion of $\pi$. We then define $\sin(x)=\cos(x-\pi/2)$ and $\tan(x)=\sin(x)/\cos(x)$.
Proof of validity is similar to the proof for $e^{x}$.
4.4 Inverse Trigonometric Functions
Finally we consider
$\frac{\arcsin(x)}{x}=\sum_{n=0}^{\infty}\frac{(2n)!}{4^{n}(n!)^{2}(2n+1)}x^{2n}=1+\frac{x^{2}}{6}+\frac{3x^{4}}{40}+\cdots$
with $|x|\leq 1$. Let $t_{n}=\frac{(2n)!}{4^{n}(n!)^{2}(2n+1)}$ and
$M_{n,w}(y)=1+\frac{t_{n}}{t_{n-1}}wy=1+\frac{(2n-1)^{2}}{2n(2n+1)}wy\ .$
Define $w=x^{2}$ and
$a_{n}(w)=M_{n,w}(M_{n+1,w}(\cdots))=1+\frac{t_{n}}{t_{n-1}}w+\frac{t_{n+1}}{t_{n-1}}w^{2}+\cdots$
so $\arcsin(x)=xa_{1}(w)$. Since $0\leq w\leq 1$, and $\{t_{n}\}$ is strictly decreasing, we can obtain the initial interval of $a_{n}(w)$ from
$1\leq a_{n}(w)\leq 1+\frac{(\arcsin(1)-1)w}{t_{n-1}}<1+\frac{4w}{7t_{n-1}}\ .$
For $n=1$ we have $t_{0}=1$ and can immediately produce 1 as the first term. Proof of validity is similar to the proof for logarithm.
4.5 Spigot Algorithms
Expressions like (7), representing an infinite sum as an infinitely nested sequence of rational functions, are central to the spigot algorithm[3, 11] for computing digits of $\pi$. One version of that algorithm uses Gosper’s “accelerated” series [4]:
$\pi=3+\frac{1\times 1}{3\times 4\times 5}\big{(}8+\frac{2\times 3}{3\times 7\times 8}\big{(}\cdots 5i-2+\frac{i(2i-1)}{3(3i+1)(3i+2)}\cdots\big{)}\big{)}\ .$
We can define
$M_{i}(x)=\cdots 5i-2+\frac{i(2i-1)}{3(3i+1)(3i+2)}x$
and $p_{i}=\lim_{i\to\infty}M_{i}(M_{i+1}(\cdots))$; then $\pi=p_{1}$. The spigot algorithm uses the fact that
$\frac{27i-12}{5}\leq p_{i}\leq\frac{27}{5}i-\frac{216}{125}\ .$ (12)
Following the same pattern as used for transcendental functions, we could obtain the CF terms of $\pi=p_{1}$ using $p_{i}=M_{n}(p_{i+1})$, taking (12) as the first interval of $p_{i}$. The spigot algorithm uses the same equation to produce digits; when the floor of the current rational expression is known to be $k$, we transform it as $M\leftarrow 10(M-k)$ to obtain an expression whose floor is the next digit.
In fact, in the same way, we could produce base-$b$ digits of $e^{x}$ (with $x$ given as a CF!) by redefining $\mbox{produce}(a,M)$ as
$M(x,y)\leftarrow b(M(x,y)-a)$
when we compute $y_{1}(x)$; after such a transformation, $M(x,y)$ represents a real number whose floor is the next digit, rather than the next CF term. For that matter, if $x$ were given as a sequence of digits, we could read it by redefining $\mbox{ingest\_x}(a,M)$ as the transformation
$x\leftarrow a+\frac{x}{b}\ .$
After this transformation, $M(x,y)$ represents the same value, with $x=[a,x_{1},x_{2}\cdots]$ replaced by $x=[x_{1},x_{2}\cdots]$; the upper and lower bounds on $M(x,y)$ would be determined by $0\leq x<b$ instead of $1<x\leq\infty$.
4.6 Related Work
Davison [2] gives an algorithm for the CF terms of $e^{\alpha}$ with $\alpha$ rational; it computes convergents using the fact that if
$\begin{pmatrix}p_{n}&q_{n}\\ r_{n}&s_{n}\end{pmatrix}=\prod_{k=0}^{n}\begin{pmatrix}(2k+1)m+\ell&(2k+1)m\\ (2k+1)m&(2k+1)m-\ell\end{pmatrix}$
then $p_{n}/r_{n}$ and $q_{n}/s_{n}$ approach $e^{\ell/m}$.
We note that standard computational software packages such as Mathematica999wolfram.com and Pari [^10] https://pari.math.u-bordeaux.fr [^11] https://wims.univ-cotedazur.fr/ wims/wims.cgi can compute CF expansions of complex expressions to high accuracy very much faster than our Haskell code. Our algorithms, however, can operate on real numbers defined by arbitrary sequences of CF terms $a_{0},a_{1},\cdots]$; more generally, on any number defined by converging sequences of rational upper and lower bounds. The algorithms could also be adapted to other functions defined by Taylor series, if we can obtain appropriate bounds in the same manner as (9) and (10), along with corresponding proofs of convergence.
5 Optimizing Computation of $e^{x}$ and $\log(x)$
In section (4.1.1), the larger $x$ is, the larger $n$ will have to be before $y_{n}(x)$ is bounded in a given interval. In the algorithm, this corresponds to recursing more deeply into $M_{(1,x)}(M_{(2,x)}(\cdots))$ before we get tight bounds on $y_{1}(x)$. Thus our Haskell implementation computes $y_{1}(x)$ only for $0\leq x<1$; when $x>1$ we use
$e^{x}=e^{\left\lfloor x\right\rfloor}e^{x-\left\lfloor x\right\rfloor}\ .$
Ordinary multiplication can be used to obtain $e^{\left\lfloor x\right\rfloor}$, since
$e=[2,1,2,1,1,4,1,1,6,1,1,8,1,1,10,\cdots]\ .$
Indeed, when the fractional part of $x$ is greater than $\frac{1}{2}$ we could use
$\sqrt{e}=[1,1,1,1,5,1,1,9,1,1,13\cdots]$
and $e^{x}=e^{\left\lfloor x\right\rfloor}e^{x-\left\lfloor x\right\rfloor-\frac{1}{2}}\sqrt{e}$.
Similarly, computation of $g(w)$ recurses less deeply when $w$ is small, i.e. when $x$ is close to 1; for $x>e$ our Haskell implementation uses $\log(x)=1+\log(x/e)$.